# UniBERT: Adversarial Training for Language-Universal Representations

### v2

@@ 3.3.2 Adversarial Training
In UniBERT, adversarial training [38] involves training a language classifier that learns to distinguish between different language inputs, using the average of the embeddings $\bar{e_{i}}$ generated by the student model. To encourage language invariance, a loss reversal technique is applied, which involves simply reversing the sign of the loss $\mathcal{L}_{\text{disc}}$ provided by the language discriminator, as shown in Equation 3:
@@ 3.3.3 Knowledge Distillation
Knowledge distillation [12] is a model compression technique in which a smaller model (known as the student) learns from a larger, pre-trained model (known as the teacher) by approximating its predictions. In UniBERT, knowledge distillation is implemented by using mBERT as the teacher model, allowing UniBERT to inherit multilingual capabilities from mBERT while maintaining a more compact and efficient structure. The objective of knowledge distillation is to minimize the Kullback-Leibler (KL) divergence between the teacher and student model outputs, as shown in Equation 4:
@@ 3.4 Training Hyperparameters
The UniBERT model was pre-trained for a total of 1M steps with a batch size of 128, utilizing a learning rate of 5e-4 with a linear scheduling approach and an initial warmup period spanning the first 50,000 steps. We selected the AdamW optimizer [39] to optimize the stability of the training. In addition, gradient clipping [40] was applied with a maximum gradient norm set to 5, effectively preventing large gradient updates. A weight decay of 0.01 was used to focus on optimization and improve the primary parameters without penalizing their magnitudes over time.
We set the training objective coefficients for the MLM loss $\lambda_{\text{MLM}}$ to 0.5, for the knowledge distillation loss $\lambda_{\text{KD}}$ to 0.4, and for the adversarial loss $\lambda_{\text{adv}}$ to 0.1. Furthermore, to refine UniBERT’s language understanding for the MLM objective, we randomly selected 15% of tokens for masking, the probability being applied uniformly, ensuring an even distribution between tokens. The MLM loss incorporated label smoothing, set to a factor of 0.7, which helped alleviate overconfidence in predictions by distributing some probability mass across non-target classes. Finally, a softmax temperature of 2 was utilized to adjust the output distribution’s sharpness, effectively balancing the confidence of the UniBERT models across potential predictions.
@@ 4 Evaluation Datasets
  • Named Entity Recognition (NER), whose goal is to identify and categorize specific entities within the text, such as people, organizations, and locations. We evaluated UniBERT on the CoNLL-2002 [41] (i.e., for the Spanish and Dutch languages) and CoNLL-2003 [42] (i.e., the English language [^5] The German subset is not available for public usage, so we did not evaluate the performance of the UniBERT models on this language.) datasets, which contain entities labeled as PER (person), ORG (organization), LOC (location), and MISC (miscellaneous). The CoNLL-2002 dataset contains approximately 12K sentences in Spanish and 24K sentences in Dutch, while CoNLL-2003 includes around 21K sentences in English. Performance was measured using the F1-score, which balances precision and recall for entity categorization.
  • Natural Language Inference (NLI), whose goal is to assess the model’s ability to understand relationships between pairs of sentences, classifying each as entailment, contradiction, or neutral. We used the Cross-lingual Natural Language Inference (XNLI) dataset [43], which includes about 400K sentence pairs for each of the 15 available languages. We used the F1-score as the primary evaluation metric, which balances precision and recall for sentence classification.
  • Question Answering (QA, which measures the ability of the models to provide answers based on a provided context. We evaluated UniBERT models on the MultiLingual Question Answering (MLQA) dataset [44], which consists of more than 90K context-question-answer triplets in seven languages. Each input includes a context paragraph and a question, with the goal of identifying the correct text span as the answer. We measured performance using F1-scores for partial matches.
  • Semantic Textual Similarity (STS), which measures the degree of semantic similarity between sentence pairs, rated on a scale from 0 (no similarity) to 5 (high similarity). We used the SemEval-2022 Task 8 (STS22) dataset [45], which provides over 9K sentence pairs with human-annotated similarity scores in ten languages. Additionally, because this dataset contains samples with the two sentences in different languages, we differentiate between fine-tuning the model on the whole dataset, which includes the samples in different languages (case called "all" in this paper) and fine-tuning the model only on the samples that contain sentences in the same language (case called "mono" in this paper). We evaluated the UniBERT performance using the Pearson correlation coefficient [46], reflecting the alignment between predicted scores and human judgments.
@@ 5.1 Named Entity Recognition
Table 2 summarizes the F1-scores achieved on the CoNLL-2002 and CoNLL-2003 datasets. The results are somewhat mixed. Among the baseline models, XLM-RoBERTa-large achieves the highest performance when individually fine-tuned, reaching F1-scores of 92.70% (en), 90.05% (es), and 93.35% (nl), for an overall average of 92.03%. In contrast, when trained jointly in all languages, its performance drops slightly (i.e., to 91.82%, 88.45%, and 91.77  for an average of 90.67%). Similarly, mBERT-base and XLM-RoBERTa-base show marginally lower scores in the "all" configuration than their "each" counterparts.
The proposed UniBERT models, while inherently more compact, display an interesting trend. For instance, UniBERT-Small achieves an average F1-score of 83.25% when individually fine-tuned, but this improves to 84.40% under joint multilingual training. A similar improvement is observed for the even smaller UniBERT-XSmall and UniBERT-XXSmall models, with average scores increasing from 77.25% to 78.80% and from 71.64% to 74.00%, respectively. These gains suggest that the joint training strategy effectively mitigates some of the capacity limitations inherent to more compact architectures, particularly for non-English languages.
@@
Distil-mBERT-base [37] each 89.26 86.50 87.02 87.59
@@
mBERT-base [2] each 90.39 88.48 90.05 89.96
@@
XLM-RoBERTa-base [17] each 91.02 88.85 91.31 90.39
@@
XLM-RoBERTa-large [17] each 92.70 90.05 93.35 92.03
@@ 5.2 Natural Language Inference
Table 3 presents the F1-scores on the XNLI dataset. The results are somewhat mixed on the NLI task, with the larger models leading overall performance. For instance, XLM-RoBERTa-large achieves an average F1-score of 82.29% when fine-tuned independently in each language, which further increases to 83.30% in joint multilingual training. Likewise, XLM-RoBERTa-base shows a notable improvement from 75.18% to 79.04% when switching from individual to joint training. In contrast, mBERT-base exhibits a moderate increase of an average of 72.64% to 73.37%. Also, Distil-mBERT-base remains relatively stable with averages of 67.68% (i.e., case "each") and 67.74% (i.e., case "all").
@@
Distil-mBERT-base [37] each 67.72 70.21 72.10 69.59 76.91 72.76 72.43 47.93 69.60 61.75 62.29 67.39 61.19 70.10 73.37 67.68
@@
mBERT-base [2] each 70.29 75.11 75.93 74.17 82.08 77.40 77.02 66.77 74.07 64.66 66.03 71.24 63.49 74.87 76.57 72.64
@@
XLM-RoBERTa-base [17] each 73.63 78.65 75.68 75.11 83.55 79.25 77.73 71.10 76.53 68.36 75.01 74.87 65.50 76.53 76.31 75.18
@@
XLM-RoBERTa-large [17] each 82.07 85.48 84.61 84.05 88.71 85.57 84.66 79.61 83.01 76.36 81.22 81.49 72.38 82.80 82.45 82.29
@@ 5.3 Question Answering
Table 4 reports the F1-scores on the MLQA dataset for each model. The baseline models consistently outperform in absolute terms. For example, XLM-RoBERTa-large achieves an average F1-score of 63.51% when individually fine-tuned, which further improves to 64.85% under joint training. Similarly, XLM-RoBERTa-base records average scores of 58.93% (i.e., case "each") and 59.97% (i.e., case "all"), mBERT-base reaches 57.86% (i.e., case "each") and 59.21% (i.e., case "all"), while Distil-mBERT-base achieves 51.95%(i.e., case "each") and 55.21% (i.e., case "all").
Our proposed UniBERT models, despite their reduced parameter count, benefit noticeably from joint multilingual fine-tuning. UniBERT-Small, for example, sees its average F1-score rise from 42.87% to 44.59%. The improvements are even more pronounced for smaller variants: UniBERT-XSmall increases from 28.43% to 36.53%, and UniBERT-XXSmall from 17.57% to 21.33%. These enhancements, which range from approximately 2% to more than 8%, highlight the advantages of using cross-lingual data during training, particularly for models with limited capacity.
@@
Distil-mBERT-base [37] each 43.28 54.32 73.42 59.68 48.90 56.07 28.04 51.95
@@
mBERT-base [2] each 51.43 59.41 78.26 65.71 56.03 63.42 30.82 57.86
@@
XLM-RoBERTa-base [17] each 52.09 59.48 77.38 64.75 61.48 65.75 31.59 58.93
@@
XLM-RoBERTa-large [17] each 58.02 65.61 80.65 70.46 68.02 69.62 32.24 63.51
@@ 5.4 Semantic Textual Similarity
Among the baseline models, mBERT-base achieves an average coefficient of 0.674 when fine-tuned on each language individually. Its performance improves to 0.723 with monolingual training and settles to 0.710 when including cross-lingual data. Similarly, XLM-RoBERTa-large obtains average scores of 0.628, 0.737, and 0.743 in the "each", "mono", and "all" settings, respectively, suggesting that joint training, especially when leveraging cross-lingual samples, can boost semantic similarity performance.
@@
Distil-mBERT-base [37] each 0.521 0.578 0.719 0.681 0.592 - 0.461 - 0.569 - 0.589
@@
mBERT-base [2] each 0.710 0.668 0.729 0.741 0.641 - 0.572 - 0.655 - 0.674
@@
XLM-RoBERTa-base [17] each 0.582 0.602 0.732 0.713 0.658 - 0.455 - 0.639 - 0.626
@@
XLM-RoBERTa-large [17] each 0.592 0.716 0.743 0.704 0.506 - 0.577 - 0.558 - 0.628

### v3

@@ 3.3.2 Adversarial Training
In UniBERT, adversarial training chen2018adversarial involves training a language classifier that learns to distinguish between different language inputs, using the average of the embeddings $\bar{e_{i}}$ generated by the student model. To encourage language invariance, a loss reversal technique is applied, which involves simply reversing the sign of the loss $\mathcal{L}_{\text{disc}}$ provided by the language discriminator, as shown in Equation 3:
@@ 3.3.3 Knowledge Distillation
Knowledge distillation hinton2015distilling is a model compression technique in which a smaller model (known as the student) learns from a larger, pre-trained model (known as the teacher) by approximating its predictions. In UniBERT, knowledge distillation is implemented by using mBERT as the teacher model, allowing UniBERT to inherit multilingual capabilities from mBERT while maintaining a more compact and efficient structure. The objective of knowledge distillation is to minimize the Kullback-Leibler (KL) divergence between the teacher and student model outputs, as shown in Equation 4:
@@ 3.4 Training Hyperparameters
The UniBERT models were pre-trained for a total of 1M steps with a batch size of 128, utilizing a learning rate of 5e-4 with a linear scheduling approach and an initial warmup period spanning the first 50,000 steps. We selected the AdamW optimizer loshchilov2018decoupled to optimize the stability of the training. In addition, gradient clipping pmlr-v28-pascanu13 was applied with a maximum gradient norm set to 5, effectively preventing large gradient updates. A weight decay of 0.01 was used to focus on optimization and improve the primary parameters without penalizing their magnitudes over time.
We set the training objective coefficients for the MLM loss $\lambda_{\text{MLM}}$ to 0.5, for the knowledge distillation loss $\lambda_{\text{KD}}$ to 0.4, and for the adversarial loss $\lambda_{\text{adv}}$ to 0.1. Furthermore, to refine UniBERT’s language understanding for the MLM objective, we randomly selected 15% of tokens for masking, the probability being applied uniformly, ensuring an even distribution between tokens. The MLM loss incorporated label smoothing, set to a factor of 0.7, which helped alleviate overconfidence in predictions by distributing some probability mass across non-target classes. Finally, a softmax temperature of 2 was utilized to adjust the output distribution’s sharpness, effectively balancing the confidence of the UniBERT models across potential predictions. We trained each UniBERT model on two NVIDIA A100 with 40GB of GPU memory and 16 CPU cores for 30 days.
@@ 4 Evaluation Datasets
  • Named Entity Recognition (NER), whose goal is to identify and categorize specific entities within the text, such as people, organizations, and locations. We evaluated UniBERT on the CoNLL-2002 tjong-kim-sang-2002-introduction (i.e., for the Spanish and Dutch languages) and CoNLL-2003 tjong-kim-sang-de-meulder-2003-introduction (i.e., the English language [^5] The German subset is not available for public usage, so we did not evaluate the performance of the UniBERT models on this language.) datasets, which contain entities labeled as PER (person), ORG (organization), LOC (location), and MISC (miscellaneous). The CoNLL-2002 dataset contains approximately 12K sentences in Spanish and 24K sentences in Dutch, while CoNLL-2003 includes around 21K sentences in English. Performance was measured using the F1-score, which balances precision and recall for entity categorization.
  • Natural Language Inference (NLI), whose goal is to assess the model’s ability to understand relationships between pairs of sentences, classifying each as entailment, contradiction, or neutral. We used the Cross-lingual Natural Language Inference (XNLI) dataset conneau2018xnli , which includes about 400K sentence pairs for each of the 15 available languages. We used the F1-score as the primary evaluation metric, which balances precision and recall for sentence classification.
  • Question Answering (QA, which measures the ability of the models to provide answers based on a provided context. We evaluated UniBERT models on the MultiLingual Question Answering (MLQA) dataset lewis2020mlqa , which consists of more than 90K context-question-answer triplets in seven languages. Each input includes a context paragraph and a question, with the goal of identifying the correct text span as the answer. We measured performance using F1-scores for partial matches.
  • Semantic Textual Similarity (STS), which measures the degree of semantic similarity between sentence pairs, rated on a scale from 0 (no similarity) to 5 (high similarity). We used the SemEval-2022 Task 8 (STS22) dataset chen2022semeval , which provides over 9K sentence pairs with human-annotated similarity scores in ten languages. Additionally, because this dataset contains samples with the two sentences in different languages, we differentiate between fine-tuning the model on the whole dataset, which includes the samples in different languages (case called "all" in this paper) and fine-tuning the model only on the samples that contain sentences in the same language (case called "mono" in this paper). We evaluated UniBERT performance using the Pearson correlation coefficient Rodgers01021988 , reflecting the alignment between predicted scores and human judgments.
@@ 5.1 Named Entity Recognition
Table 2 summarizes the F1-scores achieved on the CoNLL-2002 and CoNLL-2003 datasets. The results are somewhat mixed. Among the baseline models, XLM-RoBERTa-large achieves the highest performance when individually fine-tuned, reaching F1-scores of 92.70% (en), 90.05% (es), and 93.35% (nl), for an overall average of 92.03%. In contrast, when trained jointly in all languages, its performance drops slightly (i.e., to 91.82%, 88.45%, and 91.77%, for an average of 90.67%). Similarly, mBERT-base, XLM-RoBERTa-base, and XLM-V-base show marginally lower scores in the "all" configuration than their "each" counterparts, with the latter obtaining the highest F1-score in the "all" training configuration.
The proposed UniBERT models, while inherently more compact, display an interesting trend. For instance, UniBERT-Small achieves an average F1-score of 83.25% when individually fine-tuned, but this improves to 84.40% under multilingual joint training. A similar improvement is observed for the even smaller UniBERT-XSmall and UniBERT-XXSmall models, with average scores increasing from 77.25% to 78.80% and from 71.64% to 74.00%, respectively. These gains suggest that the joint training strategy effectively mitigates some of the capacity limitations inherent to more compact architectures, particularly for non-English languages.
@@
Distil-mBERT-base Sanh2019DistilBERTAD each 89.26 86.50 87.02 87.59
@@
mBERT-base devlin-etal-2019-bert each 90.39 88.48 90.05 89.96
@@
XLM-RoBERTa-base conneau2019cross each 91.02 88.85 91.31 90.39
@@
XLM-RoBERTa-large conneau2019cross each 92.70 90.05 93.35 92.03
@@ 5.2 Natural Language Inference
Table 3 presents the F1-scores on the XNLI dataset. The results are somewhat mixed on the NLI task, with the larger models leading overall performance. For instance, XLM-RoBERTa-large achieves an average F1-score of 82.29% when fine-tuned independently in each language, which further increases to 83.30% in multilingual joint training. Likewise, XLM-RoBERTa-base shows a notable improvement from 75.18% to 79.04% when switching from individual to joint training. In contrast, mBERT-base exhibits a moderate increase of an average of 72.64% to 73.37%. Also, Distil-mBERT-base remains relatively stable with averages of 67.68% (i.e., case "each") and 67.74% (i.e., case "all"). Finally, XLM-V-base obtained the highest performance in the "all" training configuration with an F1-score of 83.92%, its "each" training performance being slightly lower than XLM-RoBERTa-large (i.e., an F1-score of 81.92%).
@@
Distil-mBERT-base Sanh2019DistilBERTAD each 67.72 70.21 72.10 69.59 76.91 72.76 72.43 47.93 69.60 61.75 62.29 67.39 61.19 70.10 73.37 67.68
@@
mBERT-base devlin-etal-2019-bert each 70.29 75.11 75.93 74.17 82.08 77.40 77.02 66.77 74.07 64.66 66.03 71.24 63.49 74.87 76.57 72.64
@@
XLM-RoBERTa-base conneau2019cross each 73.63 78.65 75.68 75.11 83.55 79.25 77.73 71.10 76.53 68.36 75.01 74.87 65.50 76.53 76.31 75.18
@@
XLM-RoBERTa-large conneau2019cross each 82.07 85.48 84.61 84.05 88.71 85.57 84.66 79.61 83.01 76.36 81.22 81.49 72.38 82.80 82.45 82.29
@@ 5.3 Question Answering
Table 4 reports the F1-scores on the MLQA dataset for each model. The baseline models consistently outperform in absolute terms. For example, XLM-RoBERTa-large achieves an average F1-score of 63.51% when individually fine-tuned, which further improves to 64.85% under joint training. Similarly, XLM-RoBERTa-base records average scores of 58.93% (i.e., case "each") and 59.97% (i.e., case "all"), mBERT-base reaches 57.86% (i.e., case "each") and 59.21% (i.e., case "all"), while Distil-mBERT-base achieves 51.95%(i.e., case "each") and 55.21% (i.e., case "all"). XLM-V-base performs comparably to XLM-RoBERTa-large, with scores of 63.34% (i.e., case "each") and 65.57% (i.e., case "all"), showing the strongest improvement among larger models on this task.
Our proposed UniBERT models, despite their reduced parameter count, benefit noticeably from multilingual joint fine-tuning. UniBERT-Small, for example, sees its average F1-score rise from 42.87% to 44.59%. The improvements are even more pronounced for smaller variants: UniBERT-XSmall increases from 28.43% to 36.53%, and UniBERT-XXSmall from 17.57% to 21.33%. These enhancements, which range from approximately 2% to more than 8%, highlight the advantages of using cross-lingual data during training, particularly for models with limited capacity.
@@
Distil-mBERT-base Sanh2019DistilBERTAD each 43.28 54.32 73.42 59.68 48.90 56.07 28.04 51.95
@@
mBERT-base devlin-etal-2019-bert each 51.43 59.41 78.26 65.71 56.03 63.42 30.82 57.86
@@
XLM-RoBERTa-base conneau2019cross each 52.09 59.48 77.38 64.75 61.48 65.75 31.59 58.93
@@
XLM-RoBERTa-large conneau2019cross each 58.02 65.61 80.65 70.46 68.02 69.62 32.24 63.51
@@ 5.4 Semantic Textual Similarity
Among the baseline models, mBERT-base achieves an average coefficient of 0.674 when fine-tuned on each language individually. Its performance improves to 0.723 with monolingual training and settles to 0.710 when including cross-lingual data. Similarly, XLM-RoBERTa-large obtains average scores of 0.628, 0.737, and 0.743 in the "each", "mono", and "all" settings, respectively, suggesting that joint training, especially when leveraging cross-lingual samples, can boost semantic similarity performance. XLM-V-base follows a similar pattern with scores of 0.621, 0.729, and 0.730 in the respective settings.
@@
Distil-mBERT-base Sanh2019DistilBERTAD each 0.521 0.578 0.719 0.681 0.592 - 0.461 - 0.569 - 0.589
@@
mBERT-base devlin-etal-2019-bert each 0.710 0.668 0.729 0.741 0.641 - 0.572 - 0.655 - 0.674
@@
XLM-RoBERTa-base conneau2019cross each 0.582 0.602 0.732 0.713 0.658 - 0.455 - 0.639 - 0.626
@@
XLM-RoBERTa-large conneau2019cross each 0.592 0.716 0.743 0.704 0.506 - 0.577 - 0.558 - 0.628
@@
5.5 Out-of-Domain Analysis
We further performed an out-of-domain analysis of the UniBERT models for the legal domain using the MAPA dataset DeGibertBonet2022 . MAPA is a NER dataset designed for anonymization tasks in legal documents across 22 European languages. The dataset contains annotated entities that require anonymization in official documents, such as PERSON, ORGANIZATION, LOCATION, and other domain-specific categories. This evaluation allows us to assess how well our models generalize to specialized domains that differ significantly from the general Wikipedia text used during pre-training.
Table 6 presents the F1-scores achieved on the MAPA dataset. The results reveal several interesting patterns regarding the adaptability of multilingual models to domain-specific tasks. Among the baseline models, XLM-RoBERTa-base shows the highest overall performance in the "all" configuration with an average F1-score of 73.04%, followed by mBERT-base at 70.77% and Distil-mBERT-base at 70.55%. Interestingly, XLM-RoBERTa-large, despite its larger capacity, achieves a slightly lower average score of 69.51% in the joint training scenario, suggesting that model size alone does not guarantee better domain adaptation.
The most notable finding is the significant improvement that all models, especially our UniBERT variants, experience when transitioning from individual language fine-tuning to multilingual joint training. For instance, UniBERT-Small improves from an average F1-score of 46.78% in the "each" setting to 64.63% in the "all" setting, representing a relative improvement of 38.16%. Similarly, UniBERT-XSmall and UniBERT-XXSmall show substantial gains, increasing from 38.18% to 57.79% (51.36% relative improvement) and from 30.21% to 47.04% (55.71% relative improvement), respectively.
This pronounced improvement in the legal domain suggests that multilingual joint training and adversarial objectives are particularly beneficial for domain adaptation tasks. By learning language-invariant representations during pre-training, the models develop a more robust understanding of entity structures that transfers well across both languages and domains. With its specialized terminology and structured document formats, the legal domain benefits significantly from the cross-lingual representations learned through our training methodology.
Table 6: Named entity recognition results on the MAPA dataset for the legal domain. We report the model’s F1-score for each language (i.e., case "each") and for all languages at the same time (i.e., case "all"), evaluated on each language, together with their average F1-score.
Model Train bg cs da de el en es et fi fr ga hu it lt lv mt nl pt ro sk sv Avg
Distil-mBERT-base Sanh2019DistilBERTAD each 67.95 60.73 58.04 62.56 70.28 55.17 69.11 47.79 60.52 54.50 54.51 64.86 88.05 58.88 57.50 39.44 63.88 58.01 65.05 52.03 68.77 60.83
all 69.97 66.80 68.81 72.20 74.65 71.65 87.99 49.78 62.45 67.06 67.39 74.46 86.00 65.92 63.14 75.70 72.60 80.33 77.80 59.83 67.13 70.55
mBERT devlin-etal-2019-bert each 73.50 67.96 68.52 67.86 70.89 60.17 77.11 54.26 66.55 67.69 58.85 74.54 78.18 66.79 62.42 59.39 69.57 75.44 65.09 61.54 69.51 67.42
all 73.85 77.22 64.26 68.45 67.16 76.77 91.82 56.15 62.98 60.35 59.26 76.89 87.97 66.13 62.96 80.21 75.06 73.35 74.59 66.51 64.33 70.77
XLM-RoBERTa-base conneau2019cross each 66.25 60.40 63.94 57.05 50.87 52.66 49.06 46.46 64.47 48.87 53.33 56.51 89.20 52.67 49.25 34.82 61.47 56.17 59.35 50.85 72.61 56.96
all 66.53 68.55 76.34 77.03 78.30 70.52 91.39 64.58 70.03 61.51 68.76 72.28 86.08 73.68 68.71 71.72 71.88 81.34 76.89 61.98 75.80 73.04
XLM-RoBERTa-large conneau2019cross each 62.67 62.48 71.31 60.03 62.33 72.67 76.92 50.87 68.72 64.05 56.66 74.70 77.25 62.45 44.55 57.51 69.00 47.14 70.11 46.92 77.86 63.62
all 66.27 64.65 69.96 73.33 70.49 72.39 88.59 50.50 61.94 58.70 62.96 73.44 90.03 71.72 67.54 77.53 67.38 70.91 75.28 60.43 65.72 69.51
XLM-V-base liang2023xlm each 65.14 62.33 56.20 58.76 55.32 68.95 76.21 49.65 61.81 53.47 50.88 65.94 79.38 53.92 58.76 62.11 63.27 61.48 66.85 57.72 59.26 61.30
all 70.97 68.51 60.74 62.18 60.51 77.84 89.44 53.98 65.38 58.39 54.30 72.60 84.81 59.24 64.14 79.74 69.89 68.90 73.61 64.05 64.04 67.77
UniBERT-small (ours) each 57.53 44.53 49.72 49.66 51.84 45.16 47.15 29.73 44.66 44.26 45.20 57.62 70.67 42.66 38.37 7.52 55.79 43.69 59.39 45.91 51.39 46.78
all 70.09 63.95 60.15 62.55 67.49 68.10 81.58 48.26 60.15 57.20 52.42 72.18 84.79 58.92 57.48 65.27 71.46 64.83 71.26 61.71 57.41 64.63
UniBERT-xsmall (ours) each 50.83 38.16 39.47 39.65 42.08 47.38 50.03 22.36 38.58 20.30 23.01 45.98 73.47 41.35 35.44 0.00 56.04 41.16 51.76 2.25 42.54 38.18
all 69.83 63.38 51.93 51.34 62.80 58.86 71.25 37.48 56.93 53.10 45.21 59.46 73.09 53.39 49.74 57.63 65.22 54.28 63.33 63.18 51.17 57.79
UniBERT-xxsmall (ours) each 39.85 34.56 36.27 32.14 37.93 14.82 15.64 17.93 32.45 14.26 19.58 38.29 66.47 30.18 27.35 10.42 46.32 31.03 41.58 12.17 35.24 30.21
all 53.95 51.07 48.03 45.89 47.34 53.10 65.71 23.94 44.07 46.08 38.75 48.88 69.35 34.06 33.94 14.94 61.93 42.97 60.25 54.31 49.35 47.04